The lenth of a rectangle is three times the width. If the length was increased by 2 in. and the width 1 in. the new perimeter would be 62 in. what is the length and width?
Answer by Nada(11)

Sorry wrong question.
I will try to do it myself , and post the answer , it takes me a while because I study Algebra in Arabic so it's bit confusing when reading English.
I'll try my best !
Here is the asnwer :
l > is the length & w > is the width
Because a perimater of a rectangle is > 2( Width + Length )
& w = ( w+1 ) & l = ( 3w+2 )
then the equation will be 62 = 2 ( (w+1) + (3w+2) )
62 = 2w+2+6w+4
62 = 8w+6
8w = 62-6 >> w = 56/8 = 7
Then to get the length :
62 = 2(7+L)
62 = 14+2L
2L = 62-14 >> L = 48/2 = 24
To check your answer :
2(7+24) = 62 ( answer is correct)
Hope I helped.